An electric bulb is rated 60W and 120V. Calculate the resistance of its filament when it is operating normally


1
Expert's answer
2021-08-02T14:25:56-0400

Gives

P=60W

V=120v

We know that

"P=\frac{V^2}{R}\\R=\frac{V^2}{P}\\put value\\R=\frac{(120)^2}{60}=240\Omega"